# Complex structures of the Gibbons-Hawking ansatz with infinite topological type

### v1

@@ 2 Existence criteria for harmonic functions via Riesz measures
3 The complex structure of $M$ under generic assumptions
This section is devoted to describing the complex structure $J$ defined in (3) on the hyperkähler manifold $M$ of infinite topological type constructed through the Gibbons-Hawking ansatz in Section 1.1. Under the generic assumption we prove that $(M,J)$ is biholomorphic to an explicit hypersurface in $\mathbb{C}^{3}$. Our approach builds on the work of LeBrun[18], but is adapted to the case where the $S^{1}$-action has countably many fixed points.
@@ 3.1 A holomorphic $\mathbb{C}^{*}$ action on $M$
$\Psi:=-{\rm{i}}(dy\wedge\omega+Vdz\wedge dx)+(dz\wedge\omega+Vdx\wedge dy)$
@@
Then $H:M^{\pm}\to\mathbb{C}$ are holomorphic principal $\mathbb{C}^{*}$-bundles. $\mathbb{C}$ is Stein manifold. Since $\mathbb{C}^{*}$ deformation-retracts to $S^{1}$, we have
@@
Under these identifications, the vector field $\xi$ is mapped to $v\frac{\partial}{\partial v}$ and the $2$-form $\Psi$ is mapped to $\frac{dv\wedge du}{v}.$ The induced $\mathbb{C}^{*}$-action is therefore
@@ 3.2 The equivalence relation
Therefore, the complex structure can be decided only by behaviors of $f$ near the punctures. ∎
@@
$\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$.
@@
Now $H:M\to\mathbb{C}$ has nondegenerate critical point at the isolated points $q_{j}=\pi^{-1}(p_{j})$ for $p_{j}=(x_{j},y_{j},z_{j})$. Indeed, we have
$dH_{p_{j}}=\iota_{\xi}\Psi|_{p_{j}}=0$
@@
$\nabla dH(v,w)=\left(\nabla\iota_{\xi}\Psi(v,w)\right)=\left(\nabla_{v}\Psi(\xi,w)\right)-\Psi(\xi,\nabla_{v}w)=\Psi(\nabla_{v}\xi,w)$
@@
Where no confusion can arise, we drop the index $j$ for brevity. On $(U_{j},\varphi^{j})$, we have
$\Psi=\psi(w_{1},w_{2})dw_{1}\wedge dw_{2}$
@@
$\iota_{\xi}\Psi=dH=w_{1}dw_{2}+w_{2}dw_{1}=\iota_{\xi}(\psi dw_{1}\wedge dw_{2}),$
@@
$\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2})d\theta=2\pi.$
@@
$\frac{dw_{1}}{d\lambda}=\psi^{-1}w_{1}$
$\frac{dw_{2}}{d\lambda}=-\psi^{-1}w_{2}$
@@
$\oint d\lambda=2\pi{\rm{i}}=\oint\psi\frac{dw_{1}}{w_{1}}={\rm{i}}\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2})d\theta.$
@@
And the equivalence relation is decided by
@@
We now relate this local description to the Weierstrass product introduced in Section 1.2. Recall that under the generic assumption we have a countable set $\{a_{j}\}$ of pairwise distinct points of $\mathbb{C}$ without accumulation, and we define
@@
where $\delta=0$ if $a_{j}\neq 0$ for all $j$ and $\delta=1$ if $0\in\{a_{j}\}$, and where $E_{j}$ are the standard Weierstrass primary factors. Then $P(u)$ is an entire function with simple zeros exactly at the points $a_{j}$.
@@
We now give the proof of our second main result.
Proof of Theorem 1.3.
@@
$(u_{1},u_{2},u_{3})\mapsto\left\{\begin{aligned} (u_{3},u_{2})_{+}\ &\text{ if }u_{2}\neq 0,\\ \left(u_{3},\frac{1}{u_{1}}\right)_{-}\ &\text{ if }u_{1}\neq 0,\end{aligned}\right$.
which gives a holomorphic inverse to $\chi$. Hence $\chi$ is a biholomorphism from $M_{0}$ to $\Sigma\setminus\sigma$ and locally maps $U_{j}\setminus\{q_{j}\}$ to the intersection of $\Sigma\setminus\sigma$ and a neighbourhood of $(0,0,y_{j}+{\rm{i}}z_{j})$ in $\Sigma$.
By the following Hartogs’s theorem [15, Theorem 2.3.2], $\chi$ extends uniquely to a biholomorphism from $M$ to $\Sigma$. ∎
@@
Theorem 1.3 recovers, as a special case, the example of Anderson-Kronheimer-LeBrun [1] in which $\{p_{j}\}_{j}=\{(0,0,z_{j})\in\mathbb{R}^{3}\}_{j}$ with $\sum_{j=1}^{\infty}\frac{1}{|z_{j}|}<\infty$.
In this situation one has $a_{j}={\rm{i}}z_{j}$ and the Weierstrass product reduces to
@@
So far we have worked with the complex structure $J=J_{x}:dx\mapsto V^{-1}(dt+\theta),\ dy\mapsto dz$, for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to the $x$-axis. Recall we have
$J_{y}:dy\mapsto V^{-1}(dt+\theta),\ dz\mapsto dx,$
$J_{z}:dz\mapsto V^{-1}(dt+\theta),\ dx\mapsto dy.$
@@
Then for each direction $v=(v_{1},v_{2},v_{3})\in S^{2}\subset\mathbb{R}^{3}$, we obtain a compatible complex structure
@@
for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. The projection to the plane $v^{\perp}$
$\Pi_{v}=\text{Id}-vv^{T}:\mathbb{R}^{3}\to v^{\perp}\cong\mathbb{C}.$
plays the role of $H$ in this setting.
We now show that for almost such directions $v$ the previous construction applies.
Corollary 3.7.
For $v\in S^{2}$, define
$a_{j}(v):=\Pi_{v}(p_{j})\text{ and }P_{v}(u):=u^{\delta(v)}\prod_{a_{j}(v)\neq 0}E_{j}\left(\frac{u}{a_{j}(v)}\right).$
where $\delta(v)=0$ if $a_{j}(v)\neq 0$ for all $j$ and $\delta(v)=1$ if otherwise. Then for almost every $v\in S^{2}$, the points $\{a_{j}(v)\}$ are pairwise distinct, have no accumulation point in $\mathbb{C}$, and the conclusion of Theorem 1.3 holds with $H$ replaced by $\Pi_{v}$.
In particular, for such $v$, $(M,J_{v})$ is biholomorphic to the hypersurface
$u_{1}u_{2}=P_{v}(u_{3})\subset\mathbb{C}^{3}.$
Proof.
@@
By the Borel-Cantelli lemma, for $A_{n}:=\{v\in S^{2}:\ v\in A_{j,n}\text{ for infinitely many }j\}=\limsup_{j\to\infty}A_{j,n}$,
@@
is a countable set. Hence for every $v\in S^{2}\setminus B$, $\Pi_{v}$ is injective on $\{p_{j}\}$. Then Theorem 1.3 completes the proof.
@@ 4 Non-injective projection and the minimal resolution
While the previous discussion focused on the case where $H$ is bijective from $\{p_{j}\}_{j}$ to their images, where the complex structure of the infinite topological type hyperkähler manifold $M$ is realized by a hypersurface in $\mathbb{C}^{3}$, the aim of this section is to relax this injectivity assumption and to analyse the resulting non-generic configurations. When $H$ fails to be one-to-one, $M$ should then be interpreted not as the hypersurface itself, but as its minimal resolution of singularities.
@@
Without loss of generality, assume that there are only two fixed points under the $S^{1}$ action: $q_{1}=\pi^{-1}(p_{1}),\ q_{2}=\pi^{-1}(p_{2})$ where $p_{1}=(0,0,0)$ and $p_{2}=(1,0,0)$. In this situation, Chen-Chen [4] has shown that the resulting hyperkähler manifold $M$ is biholomorphic to the minimal resolution of the hypersurface of $\mathbb{C}^{3}$ defined by $u_{1}u_{2}=u_{3}^{2}$.
@@
and let $Bl_{0}(S)$ denoted the blowing-up of $S$ at point $(0,0,0)\in\mathbb{C}^{3}$:
$Bl_{0}(S)$ $=\{((u_{1},u_{2},u_{3}),[U_{1},U_{2},U_{3}]):u_{1}U_{2}=u_{2}U_{1},\ u_{1}U_{3}=u_{3}U_{1},\ u_{2}U_{3}=u_{3}U_{2},\ U_{1}U_{2}=U_{3}^{2}\}$
@@
We now generalize this picture to the case of countably many monopole points, where several monopoles may project to the same point under $H$.
Let $\{p_{j}\}_{j\in\mathbb{N}^{+}}\subset\mathbb{R}^{3}$ be a countable set of distinct points such that
@@
Write $a_{j}=y_{j}+{\rm{i}}z_{j}=H(p_{j})$, and let $\{b_{1},...,b_{n},$...\} denote the set of distinct values in $\{a_{j}\}$. We assume that $\{b_{k}\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$ and for all $k\geq 1$,
@@
Relabel the monopoles as $\{p_{k,l}\}_{k,l}$ so that
$H(p_{k,l})\equiv b_{k},\ p_{k,l}=(x_{k,l},y_{k,l},z_{k,l}).$
Define an entire function
$P(u):=u^{m_{0}}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{m_{k}},$
where $c_{k}$ are properly chosen to ensure the infinite product to converge uniformly on every compact subset of $\mathbb{C}$. Let $S\subset\mathbb{C}^{3}$ be the hypersurface defined by
@@
$f_{\alpha\beta}(u)=u^{J_{0}(\alpha)-J_{0}(\beta)}\prod_{k:b_{k}\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}}\right)\right)^{J_{k}(\alpha)-J_{k}(\beta)}.$
Proof of Theorem 1.4.
@@
and the restriction $\chi|_{M\setminus E}:M\setminus E\to S\setminus\{(0,0,b_{k})\}_{k=1}^{n}$ is a biholomorphism. Since $\chi$ is proper, it follows that $\chi:M\to S$ is a resolution of singularities of $S$.
@@
Here we have $g(E_{k,l})=0$ and $K_{M}\cdot E_{k,l}=0$ since $M$ is a hyperkähler manifold and $K_{M}$ is trivial. Hence $E_{k,l}^{2}=-2$ for all $k,l$. Thus every component of the exceptional divisor has self-intersection 2-2, and no component over any $b_{k}$ can be further contracted. Therefore $\chi:M\to S$ is a minimal resolution. ∎
Finally, recall that for a direction $v=(v_{1},v_{2},v_{3})\in S^{2}\subset\mathbb{R}^{3}$ we could choose complex structure $J_{v}=v_{1}J_{x}+v_{2}J_{y}+v_{3}J_{z}$ so that the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. If $v\in S^{2}$ is the direction satisfying that the projected points $a_{j}(v):=\Pi_{v}(p_{j})\in\mathbb{C}$ have pairwise distinct values $\{b_{k}(v)\}_{k}$ with no accumulation point on $\mathbb{C}$, and the multiplicity at each $b_{k}(v)$ is finite:
$m_{0}(v):=\#\{j\in\mathbb{N}^{+}:a_{j}(v)=0\}<\infty,$
$m_{k}(v):=\#\{j\in\mathbb{N^{+}}:a_{j}(v)=b_{k}(v)\}<\infty.$
Then we can define
$P_{v}(u):=u^{m_{0}(v)}\prod_{k:b_{k}(v)\neq 0}\left(E_{c_{k}}\left(\frac{u}{b_{k}(v)}\right)\right)^{m_{k}(v)}$
where $c_{k}(v)$ are properly chosen. For such a direction $v$, the above argument applies verbatim to $(M,J_{v})$, and show that $(M,J_{v})$ is biholomorphic to the minimal resolution of the hypersurface in $\mathbb{C}^{3}$ defined by $u_{1}u_{2}=P_{v}(u_{3})$, whose singularity type is $(A_{m_{1}-1},\cdots,A_{m_{k}-1},\cdots)$.

### v2

@@ 2 Existence criteria for harmonic functions via Riesz measures
3 Hypersurface realizations in the hyperkähler family
In this section we describe the complex structures in the hyperkähler family of $M$ which arise from certain directions. We begin by fixing a direction along the $x$-axis, and show that $(M,J_{x})$ is biholomorphic to a smooth hypersurface in $\mathbb{C}^{3}$, which is recorded in Lemma 3.1. We then prove that for almost every direction on $\mathbb{S}^{2}$ and the corresponding complex structure $J_{v}$, the same description applies, thereby obtaining Theorem 1.3. Our approach builds on the work of LeBrun [19], but is adapted to the case where the $S^{1}$-action has countably many fixed points.
Let $H:\mathbb{R}^{3}\to\mathbb{C}$ be the projection $H(x,y,z):=y+{\rm{i}}z$ and $a_{j}:=H(p_{j})=y_{j}+{\rm{i}}z_{j}$. We say that the configuration $\{p_{j}\}_{j=1}^{\infty}$ is generic if the projection
$H:\{p_{j}\}_{j=1}^{\infty}\to\mathbb{C}$
is injective, that is, $\{a_{j}\}_{j=1}^{\infty}$ are pairwise distinct. Note that the projection $H$ here is exactly $\Pi_{x}$ introduced in Section 1.2.
In the sequel we shall mainly work with the complex structure $J_{x}$ and for simplicity, we write $J:=J_{x}$. For later use it is convenient to describe it in slightly more concrete coordinates. Writing $\omega={{\mathrm{d}}}t+\theta$ for some $1$-form $\theta$ such that ${{\mathrm{d}}}\theta=\star{{\mathrm{d}}}V$, where $t$ is the angular coordinate along the $S^{1}$-fibres, the complex structure $J$ can be characterized by
${{\mathrm{d}}}x\mapsto V^{-1}({{\mathrm{d}}}t+\theta),\ {{\mathrm{d}}}y\mapsto{{\mathrm{d}}}z.$ (9)
Our key Lemma is the following description for $(M,J)$ under the generic hypothesis on the puncture points. For this we recall the standard Weierstrass factors
Lemma 3.1.
Suppose that the puncture points $\{p_{j}\}$ satisfy
$\sum_{j=2}^{\infty}\frac{1}{\|p_{1}-p_{j}\|}<\infty,$
and that the configuration $\{p_{j}\}_{j=1}^{\infty}$ is generic. If, in addition, the set $\{a_{j}\}_{j=1}^{\infty}$ has no accumulation point on $\mathbb{C}$, then
$P(u)=u^{\delta}\prod_{a_{j}\neq 0}E_{j}\left(u/a_{j}\right)$
where
$\delta=\left\{\begin{aligned} 0\ \ &\text{ if }0\notin\{a_{j}\},\\ 1\ \ &\text{ if }0\in\{a_{j}\},\end{aligned}\right$.
is an entire function. Moreover $(M,J)$ is biholomorphic to the hypersurface in $\mathbb{C}^{3}$ given by
$u_{1}u_{2}=P(u_{3}).$
@@ 3.1 A holomorphic $\mathbb{C}^{*}$ action on $M$
$\Psi:=-{\rm{i}}({{\mathrm{d}}}y\wedge\omega+V{{\mathrm{d}}}z\wedge{{\mathrm{d}}}x)+({{\mathrm{d}}}z\wedge\omega+V{{\mathrm{d}}}x\wedge{{\mathrm{d}}}y)$
@@
Then $H:M^{\pm}\to\mathbb{C}$ are holomorphic principal $\mathbb{C}^{*}$-bundles. $\mathbb{C}$ is a Stein manifold. Since $\mathbb{C}^{*}$ deformation-retracts to $S^{1}$, we have
@@
Under these identifications, the vector field $\xi$ is mapped to $v\frac{\partial}{\partial v}$ and the $2$-form $\Psi$ is mapped to $\frac{{{\mathrm{d}}}v\wedge{{\mathrm{d}}}u}{v}.$ The induced $\mathbb{C}^{*}$-action is therefore
@@ 3.2 The equivalence relation
Therefore, the complex structure can be determined only by behaviors of $f$ near the punctures. ∎
@@
and
$\delta=\left\{\begin{aligned} 0\ \ &\text{ if }0\notin\{a_{j}\},\\ 1\ \ &\text{ if }0\in\{a_{j}\}.\end{aligned}\right$.
@@
Now $H:M\to\mathbb{C}$ has nondegenerate critical points at the isolated points $q_{j}=\pi^{-1}(p_{j})$ for $p_{j}=(x_{j},y_{j},z_{j})$. Indeed, we have
${{\mathrm{d}}}H_{p_{j}}=\iota_{\xi}\Psi|_{p_{j}}=0$
@@
$\nabla{{\mathrm{d}}}H(v,w)=\left(\nabla\iota_{\xi}\Psi(v,w)\right)=\left(\nabla_{v}\Psi(\xi,w)\right)-\Psi(\xi,\nabla_{v}w)=\Psi(\nabla_{v}\xi,w)$
@@
When no confusion can arise, we drop the index $j$ for brevity. On $(U_{j},\varphi^{j})$, we have
$\Psi=\psi(w_{1},w_{2}){{\mathrm{d}}}w_{1}\wedge{{\mathrm{d}}}w_{2}$
@@
$\iota_{\xi}\Psi={{\mathrm{d}}}H=w_{1}{{\mathrm{d}}}w_{2}+w_{2}{{\mathrm{d}}}w_{1}=\iota_{\xi}(\psi{{\mathrm{d}}}w_{1}\wedge{{\mathrm{d}}}w_{2}),$
@@
$\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2}){{\mathrm{d}}}\theta=2\pi.$
@@
$\frac{{{\mathrm{d}}}w_{1}}{{{\mathrm{d}}}\lambda}=\psi^{-1}w_{1}$
$\frac{{{\mathrm{d}}}w_{2}}{{{\mathrm{d}}}\lambda}=-\psi^{-1}w_{2}$
@@
$\oint{{\mathrm{d}}}\lambda=2\pi{\rm{i}}=\oint\psi\frac{{{\mathrm{d}}}w_{1}}{w_{1}}={\rm{i}}\int_{0}^{2\pi}\psi(e^{{\rm{i}}\theta}w_{1},e^{-{\rm{i}}\theta}w_{2}){{\mathrm{d}}}\theta.$
@@
And the equivalence relation is determined by
@@
Recall that under the generic assumption we have a countable set $\{a_{j}\}$ of pairwise distinct points of $\mathbb{C}$ without accumulation, and we define
@@
where
$\delta=\left\{\begin{aligned} 0\ \ &\text{ if }0\notin\{a_{j}\},\\ 1\ \ &\text{ if }0\in\{a_{j}\},\end{aligned}\right$.
and where $E_{j}$ are the standard Weierstrass factors. Then $P(u)$ is an entire function with simple zeros exactly at the points $a_{j}$.
@@
We now complete the proof of our key Lemma.
Proof of Lemma 3.1.
@@
$(u_{1},u_{2},u_{3})\mapsto\left\{\begin{aligned} &(u_{3},u_{2})_{+}\ &\text{ if }u_{2}\neq 0,\\ &\left(u_{3},\frac{1}{u_{1}}\right)_{-}\ &\text{ if }u_{1}\neq 0,\end{aligned}\right$.
and this map gives a holomorphic inverse to $\chi$. Hence $\chi$ is a biholomorphism from $M_{0}$ to $\Sigma\setminus\sigma$ and locally maps $U_{j}\setminus\{q_{j}\}$ to the intersection of $\Sigma\setminus\sigma$ and a neighbourhood of $(0,0,y_{j}+{\rm{i}}z_{j})$ in $\Sigma$.
By the following Hartogs’s theorem [16, Theorem 2.3.2], $\chi$ extends uniquely to a biholomorphism from $M$ to $\Sigma$. This proves Lemma 3.1
@@
Lemma 3.1 recovers, as a special case, the example of Anderson-Kronheimer-LeBrun [1] in which $\{p_{j}\}_{j}=\{(0,0,z_{j})\in\mathbb{R}^{3}\}_{j}$ with $\sum_{j=1}^{\infty}\frac{1}{|z_{j}|}<\infty$.
In this situation, $a_{j}={\rm{i}}z_{j}$ and the Weierstrass product reduces to
@@
3.3 Transition to other directions in $\mathbb{S}^{2}$
So far we have worked with the complex structure $J=J_{x}:{{\mathrm{d}}}x\mapsto V^{-1}({{\mathrm{d}}}t+\theta),\ {{\mathrm{d}}}y\mapsto{{\mathrm{d}}}z$, for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to the $x$-axis. Recall that we have
$J_{y}:{{\mathrm{d}}}y\mapsto V^{-1}({{\mathrm{d}}}t+\theta),\ {{\mathrm{d}}}z\mapsto{{\mathrm{d}}}x,$
$J_{z}:{{\mathrm{d}}}z\mapsto V^{-1}({{\mathrm{d}}}t+\theta),\ {{\mathrm{d}}}x\mapsto{{\mathrm{d}}}y.$
@@ 3.3 Transition to other directions in $\mathbb{S}^{2}$
Then for each direction $v=(v_{1},v_{2},v_{3})\in\mathbb{S}^{2}\subset\mathbb{R}^{3}$, we obtain a compatible complex structure
@@
for which the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. The projection to the plane $v^{\perp}$ plays the role of $H$ in this setting.
For $v\in\mathbb{S}^{2}$, recall $a_{j}(v)=\Pi_{v}(p_{j})$ and we define the entire function
$P_{v}(u):=u^{\delta(v)}\prod_{a_{j}(v)\neq 0}E_{j}\left(\frac{u}{a_{j}(v)}\right).$
where
$\delta(v)=\left\{\begin{aligned} 0\ \ &\text{ if }0\notin\{a_{j}(v)\},\\ 1\ \ &\text{ if }0\in\{a_{j}(v)\}.\end{aligned}\right$.
We now show that for almost every directions $v\in\mathbb{S}^{2}$, the assumptions of Lemma 3.1 are satisfied and the previous construction applies.
Proof of Theorem 1.3.
@@
By the Borel-Cantelli lemma, for $A_{n}:=\{v\in\mathbb{S}^{2}:\ v\in A_{j,n}\text{ for infinitely many }j\}=\limsup_{j\to\infty}A_{j,n}$,
@@
is a countable set. Hence for every $v\in\mathbb{S}^{2}\setminus B$, $\Pi_{v}$ is injective on $\{p_{j}\}$.
Putting these facts together, we conclude that for almost every $v\in\mathbb{S}^{2}$, the points $\{a_{j}(v)\}$ are pairwise distinct and have no accumulation point in $\mathbb{C}$. For such $v$, the conclusion of Lemma 3.1 holds with $H$ replaced by $\Pi_{v}$ and $J$ by $J_{v}$. Therefore, for such $v$, $(M,J_{v})$ is biholomorphic to the hypersurface
$\{(u_{1},u_{2},u_{3})\in\mathbb{C}^{3}:u_{1}u_{2}=P_{v}(u_{3})\}\subset\mathbb{C}^{3}.$
This is exactly the statement of Theorem 1.3. ∎
@@ 4 Non-injective projection and the minimal resolution
While the previous discussion focused on the case where $\Pi_{v}$ is injective from $\{p_{j}\}_{j}$ to their images, in this section we study the case where several punctures project to the same point in $\mathbb{C}$. In this case $M$ can no longer be described directly as a hypersurface in $\mathbb{C}^{3}$. Instead, under certain conditions we show that $M$ is realized as the minimal resolution of a singular hypersurface. As in Section 3, we first treat the case of a fixed projection direction along the $x$-axis, assuming that only finitely many punctures lie over each point of the image. Then, by replacing this fixed direction with an arbitrary direction $v\in\mathbb{S}^{2}$ and working with the corresponding complex structure $J_{v}$, we obtain Theorem 1.4.
@@
Without loss of generality, assume that there are only two fixed points under the $S^{1}$ action: $q_{1}=\pi^{-1}(p_{1}),\ q_{2}=\pi^{-1}(p_{2})$ where $p_{1}=(0,0,0)$ and $p_{2}=(1,0,0)$. In this situation, Chen-Chen [4] has shown that the resulting hyperkähler manifold $M$ is biholomorphic to the minimal resolution of the hypersurface in $\mathbb{C}^{3}$ defined by $u_{1}u_{2}=u_{3}^{2}$.
@@
and let $Bl_{0}(S)$ denote the blowing-up of $S$ at point $(0,0,0)\in\mathbb{C}^{3}$:
$Bl_{0}(S)$ $=\{((u_{1},u_{2},u_{3}),[U_{1},U_{2},U_{3}]):u_{j}U_{k}=u_{k}U_{j}\ (1\leq j<k\leq 3),U_{1}U_{2}=U_{3}^{2}\}$
@@
The two-centre model already exhibits the essential phenomenon that will reappear in the general situation. In the general case, several punctures may lie over each value of the projection. We now turn to this general configuration for a fixed projection direction.
Recall that $H:\mathbb{R}^{3}\to\mathbb{C}$, $H(x,y,z)=y+{\rm{i}}z$, is the projection along $x$-axis and $a_{j}=H(p_{j})=y_{j}+{\rm{i}}z_{j}$. Let $\{b_{1},b_{2},\cdots\}$ be the set of distinct values of $\{a_{j}\}$. We shall state the main Lemma for the fixed direction along the $x$-axis and the corresponding complex structure $J=J_{x}$.
Lemma 4.2.
For countable distinct points $p_{j}\in\mathbb{R}^{3},\ j\in\mathbb{N^{+}}$, suppose that
@@
Assume that
  • (1).
    the set $\{b_{k}\}_{k=1}^{\infty}$ has no accumulation point on $\mathbb{C}$;
  • (2).
    for each $k\geq 1$, the following multiplicities are finite:
@@
Relabel the points $\{p_{j}\}_{j}$ as $\{p_{k,l}\}_{k,l}$ so that $H^{-1}(b_{k})=\{p_{k,l}:1\leq l\leq m_{k}\}$ for all $k\in\mathbb{N}^{+}$. Using suitable Weierstrass factors, we define an entire function
$P(u):=u^{m_{0}}\prod_{k:b_{k}\neq 0}\left(E_{\ell_{k}}\left(\frac{u}{b_{k}}\right)\right)^{m_{k}}$
where the positive integers $\ell_{k}$ are properly chosen so that the infinite product converges uniformly on every compact subset of $\mathbb{C}$.
Let $S\subset\mathbb{C}^{3}$ be the surface defined by
@@
Then $(M,J)$ is biholomorphic to the minimal resolution of $S$. The singularity type is $(A_{m_{1}-1},\cdots,A_{m_{k}-1},\cdots)$ for each $k$ with $m_{k}\geq 2$.
Proof.
@@
$f_{\alpha\beta}(u)=u^{J_{0}(\alpha)-J_{0}(\beta)}\prod_{k:b_{k}\neq 0}\left(E_{\ell_{k}}\left(\frac{u}{b_{k}}\right)\right)^{J_{k}(\alpha)-J_{k}(\beta)}.$
@@
and the restriction $\chi|_{M\setminus E}:M\setminus E\to S\setminus\{(0,0,b_{k})\}_{k}$ is a biholomorphism. Since $\chi$ is proper, it follows that $\chi:M\to S$ is a resolution of singularities of $S$.
@@
Here we have $g(E_{k,l})=0$ and $K_{M}\cdot E_{k,l}=0$ since $M$ is a hyperkähler manifold and $K_{M}$ is trivial. Hence $E_{k,l}^{2}=-2$ for all $k,l$. Thus every component of the exceptional divisor has self-intersection 2-2, and no component over any $b_{k}$ can be further contracted. Therefore $\chi:M\to S$ is a minimal resolution. This proves the Lemma 4.2. ∎
Finally, recall that for a direction $v=(v_{1},v_{2},v_{3})\in\mathbb{S}^{2}\subset\mathbb{R}^{3}$ we may choose complex structure
$J_{v}=v_{1}J_{x}+v_{2}J_{y}+v_{3}J_{z}$
so that the $\mathbb{R}^{3}$-projections of the $\mathbb{C}^{*}$-orbits are parallel to $v$. The projection $\Pi_{v}$ then plays the role of $H$ in this setting. For $v\in\mathbb{S}^{2}$, recall $a_{j}(v)=\Pi_{v}(p_{j})\in\mathbb{C}$ and denote the distinct values of $\{a_{j}(v)\}_{j}$ by $\{b_{k}(v)\}_{k}$ with multiplicities
$m_{0}(v):=\#\{j\in\mathbb{N}^{+}:a_{j}(v)=0\},$
$m_{k}(v):=\#\{j\in\mathbb{N^{+}}:a_{j}(v)=b_{k}(v)\}.$
Proof of Theorem 1.4.
Under the hypothesis of Theorem 1.4, $\{a_{j}(v)\}_{j}$ have no accumulation point in $\mathbb{C}$, and each multiplicity $m_{k}(v)$ is finite. We can define the corresponding entire function as
$P_{v}(u):=u^{m_{0}(v)}\prod_{k:b_{k}(v)\neq 0}\left(E_{\ell_{k}(v)}\left(\frac{u}{b_{k}(v)}\right)\right)^{m_{k}(v)}$
where $\ell_{k}(v)$ are properly chosen so that the product converges uniformly on every compact subset of $\mathbb{C}$. For such a direction $v$, the above argument in Lemma 4.2 applies verbatim to $(M,J_{v})$ with $H$ replaced by $\Pi_{v}$ and $J$ replaced by $J_{v}$. Thus $(M,J_{v})$ is biholomorphic to the minimal resolution of the hypersurface in $\mathbb{C}^{3}$ defined by
$u_{1}u_{2}=P_{v}(u_{3}).$
The singularity type over $u_{3}=b_{k}$ is of type $A_{m_{k}(v)-1}$ for each $k$ with $m_{k}(v)\geq 2$. This proves Theorem 1.4. ∎
Remark 4.3.
Hattori’s related work [13] studied hyperkähler manifolds of type $A_{\infty}$ via the hyperkähler quotient description and obtains a closely related classification. Our result is parallel to his in some sense.